# Constant-Size Certificates for Leader Election in Chordal Graphs and Related Classes

### v1

@@ 2.1 Model
2.2 Warm Up: A Certificate for Leader Election in Trees
We give a simple certificate that guarantees a unique sink in the class of trees. While not new (e.g., one could easily construct a certificate following the self–stabilizing leader election algorithm of [antonoiu1996trees]), we believe it gives a good intuition behind the certificates and proofs in the following sections. The certificate of each node consists of the orientation of edges incident to the node. A correct configuration is defined by two conditions, T1 and T2. If both T1,T2 hold at every node of a tree $T$, then $T$ has a unique sink which we define to be the leader.
@@ 2.2 Warm Up: A Certificate for Leader Election in Trees
In an unlabeled tree $T$, any configuration $\vec{T}$ satisfying T1 and T2 contains a unique sink.
Proof 2.2.
@@
3 Chordal Graphs
We generalize the certification scheme from Section 2.2 to ensure the validity of a certificate assignment, which guarantees acyclicity and a unique sink in chordal graphs. Each node’s certificate is the orientation of its incident edges. We refer to the nodes of $G$ and the certificate assignment as a configuration $\vec{G}$. A graph is chordal if every induced cycle is of length three. Every chordal graph has a simplicial vertex $v$, that is, a vertex $v$ such that $N(v)$ is a clique. We define a configuration to be correct for leader election, if for every vertex $v$ of a chordal graph, the following local conditions hold:
@@ 3 Chordal Graphs
Observe that the verification of C3 is possible, since the view by definition contains the orientation of edges between neighboring nodes.
@@
[Existence] A correct configuration with respect to C1–C3 exists and can be constructed as follows: direct all edges incident to a simplicial vertex $u$ toward $N(u)$, then iteratively restrict the graph to the subgraph of remaining unoriented edges and orient edges away from simplicial vertices.
@@
Suppose now that $u$ has no outneighbor, (i.e., $u$ is a sink of $\vec{G}$). For any neighbor $v$ of $u$, any outneighbor $w$ of $v$ in $\vec{G}\setminus\{u\}$ must be a neighbor of $u$ in $\vec{G}$ from C2. Consequently, $N^{+}(v)\subseteq N[u]$. We claim that there exists a vertex $t^{\prime}\in N(u)$ that has no outneighbor in $N(u)$. Suppose otherwise. Then, every $v\in N(u)$ has an outneighbor in $N(u)$ and $\vec{G}\setminus\{u\}$ contains a cycle, which is a contradiction by Lemma 3.1. Since $N(u)$ is a complete graph, this $t^{\prime}$ is unique. Furthermore, as we have already shown that $t^{\prime}$ has no outneighbor outside of $N[u]$, $t^{\prime}$ is the unique sink of $\vec{G}\setminus\{u\}$. Since $^{\prime}\in N^{-}(u), tt$ is not a sink in $\vec{G}$. Therefore, $u$ is the unique sink of $\vec{G}$.
Theorem 3.5 directly follows from Section˜3 and Lemmas˜3.1 and 3.3.
@@ 4 Dismantlable graphs
This section is divided into two parts: a certification scheme ensuring a unique sink in $K_{4}$-free dismantlable graphs and a certification scheme for a rooted spanning tree, given a unique root, in dismantlable graphs. Let $G$ be a graph and $u,v$ be two vertices in $V(G)$. We say that $v$ is dominated by $u$ in $G$ if $N[v]\subseteq N[u]$. A graph is dismantlable if there exists an ordering $v_{1},\ldots,v_{n}$ of $V(G)$ such that in $G[\{v_{1},\ldots,v_{i}\}]$, $v_{i}$ is dominated by some $v_{j}$ with $j<i$. So for each $i>1$, there exists $j<i$ such that $N[v_{i}]\cap\{v_{1},\ldots,v_{i}\}\subseteq N[v_{j}]$. Dismantlable graphs are exactly the class of cop-win graphs in the classic Cops and Robbers game [aigner1984cops, nowakowski1983vertex, quilliot1978jeux] and include well-structured families such as Bridged graphs and Helly graphs, which are known for their convexity and metric properties.
@@ 4.1 Leader election in $K_{4}$-free dismantlable graphs
In this section we only consider $K_{4}$-free dismantlable graphs. As in Section 3, the certificate at each node consists of an orientation on the edges incident to the node and we write $\vec{G}$ to denote a configuration. We define a locally correct configuration to be one that satisfies the following rules for every node $v$ of $G$. We show that if all rules are satisfied, then $G$ contains a unique sink that we define to be the leader.
@@
[Existence] Any dismantlable graph admits an acyclic orientation satisfying D1–D3. {claimproof} Let $v_{1},v_{2}\ldots,v_{n}$ be the dismantling order of the nodes of some $K_{4}$-free dismantlable graph $G$. Orient each edge between two nodes $v_{i}$ and $v_{j}$ as $\overrightarrow{v_{i}v_{j}}$ if and only if $i>j$ in the dismantling order. This orientation satisfies D1 and D3. For D2, consider any node $v$. If $N^{+}(v)=\emptyset$, D2 is immediately satisfied for $v$. So suppose $N^{+}(v)\neq\emptyset$. By dismantlability, every node $v$ is dominated by some later node $u$ in the dismantling order. Then $\overrightarrow{vu}$ is an edge of the orientation and for every other outneighbor $w$ of $v$, $w\in N(u)$. Hence, $N^{+}(v)\subseteq N[u]$, satisfying D2.
However, notice that in the case of $K_{4}$-free dismantlable graphs, one cannot hope to have a property similar to Lemma˜3.1. Indeed, as in [chalopin2024stabilising], one can construct configurations satisfying D1-D3 that contain directed cycles. An example is shown in Figure˜3.
Figure 3: A $K_{4}$-free dismantlable graph and a certificate assignment that satisfies D1–D3 but is not acyclic. The square node dominates all cyclic nodes. The edges forming the cycle are marked in orange.
@@
In the remainder of the proof we show that D2 is satisfied in $\vec{G}^{\prime}$ and that $\vec{G}$ contains a unique sink. We distinguish two cases, depending on whether $uv$ is directed as $\overrightarrow{uv}$ or $\overrightarrow{vu}$.
@@
We first show that D2 is satisfied in $\vec{G}^{\prime}$. Consider a vertex $w\neq v$ such that $w\in N^{-}(v)$. If $w\neq u$, since $u$ dominates $v$, $u$ must be a neighbor of $w$. Given that $N^{+}(u)\cap N^{-}(v)=\emptyset$, it follows that $w\in N^{-}(u)$. By Lemma˜4.3, we can assume $w^{+}\neq v$, thus D2 is satisfied for $w$ in $\vec{G}^{\prime}$.
We now show that D2 holds for $u$. Since $N^{+}(u)\subseteq N[v]$ and $N^{+}(u)\cap N^{-}(v)=\emptyset$, we have $N^{+}(u)\subseteq N^{+}(v)\cup\{v\}$. Given that $N^{+}(v)\subseteq N(v)\subseteq N[u]$ and $v\in N^{+}(u)$, by D3, we obtain $N^{+}(v)\subseteq N^{+}(u)$. Therefore, $N^{+}(u)=N^{+}(v)\cup\{v\}$. Hence, either $N^{+}(v)=\emptyset$ and $N^{+}(u)=\{v\}$, or there exists $v^{+}\in N^{+}(v)$ such that $N^{+}(v)\subseteq N[v^{+}]$, implying $N^{+}(u)=N^{+}(v)\cup\{v\}\subseteq N[v^{+}]$. In either case, D2 is satisfied for $u$ in $\vec{G}^{\prime}$.
By induction hypothesis $\vec{G}^{\prime}$ contains a unique sink $t$. Notice that as $N^{+}(u)=N^{+}[v]$, $v$ is a sink of $\vec{G}$ if and only if $u$ is a sink of $\vec{G}^{\prime}$. Therefore, if $t\neq u$ in $\vec{G}^{\prime}$, $v$ is not a sink in $\vec{G}$ and $t$ is the unique sink in $\vec{G}$. If $t=u$, then $t$ is not a sink in $\vec{G}$ but $v$ is, consequently $v$ is the unique sink of $\vec{G}^{\prime}$. In either case, $\vec{G}$ contains a unique sink concluding the proof of ˜2.1.1.
@@
We now show that D2 holds in $\vec{G}^{*}$. For any $\notin(N^{+}(u)\cap N^{-}(v))\cup\{u\}, xx$ has the same set of outneighbors in $\vec{G}$ and in $\vec{G}^{*}$, and thus D2 holds for $x$ in $\vec{G}^{*}$ as it holds in $\vec{G}$. Since $N^{+}(u)\subseteq N(v)$ in $\vec{G}$, since $N^{+}(u)$ in $\vec{G}^{*}$ is a subset of $N^{+}(u)$ in $\vec{G}$, and since $v\in N^{+}(u)$ in $\vec{G}^{*}$, D2 also holds for $u$ in $\vec{G}^{*}$ (with $u^{+}=v$). Consider now $w\in(N^{+}(u)\cap N^{-}(v))$ and let $w^{+}\in N^{+}(w)$ such that $N^{+}(w)\subseteq N[w^{+}]$ in $\vec{G}^{\prime}$. As $u\notin N^{+}(w)$ in $\vec{G}$, $w^{+}\neq u$. Since $v\in N^{+}(w)$, $w^{+}\in N[v]\setminus\{u\}\subseteq N(u)$. Hence, $w^{+}\in N(u)$, implying $u\in N[w^{+}]$ and thus in $\vec{G}^{*}$ we also have $N^{+}(w)\subseteq N[w^{+}]$. Therefore, D2 is satisfied at $w$ in $\vec{G}^{*}$. Hence D2 holds in $\vec{G}^{*}$.
@@
We first show that D2 is satisfied in $\vec{G}^{\prime}$. For D2 to be violated, $v$ should be the only choice for $w^{+}$, for some $w\neq u$. If $N^{+}(w)=\{v\}$, D2 trivially holds for $w$ in $\vec{G}^{\prime}$. If $u\in N^{+}(w)$, D2 holds for $w$ by Lemma˜4.3. Suppose now there exists $z\in N^{+}(w)\setminus\{v\}$. Then $z\in N(v)$ and from $u$ dominating $v$, $\{w,z\}\in N(u)$. Hence, $v,u,w,z$ induce a $K_{4}$, contradiction. So we can assume that $w^{+}\neq v$ for all $w\in V$, and D2 holds for all nodes in $\vec{G}^{\prime}$.
By induction hypothesis, $\vec{G}^{\prime}$ contains a unique sink $t$. We consider two cases, based on the orientation of $u^{+}v$.
@@
If $t$ is not a sink in $\vec{G}$, $t\in N^{-}(v)$ and thus, $t\in N(u)$. Observe that it cannot be that $t=u^{+}$ since we have assumed $\overrightarrow{vu^{+}}$, which means that $vt$ is oriented as $\overrightarrow{vt}$ and $t$ would still be a sink in $\vec{G}$ contradicting our assumption. So suppose $t\neq u^{+}$. Since $t$ is a sink in $\vec{G}^{\prime}$, $t\in N^{+}(u)$ and thus $t\in N(u^{+})$, forming a $K_{4}$ among $u,v,t,u^{+}$. Therefore, $t\notin N^{-}(v)$ and $t$ is also a sink in $\vec{G}$. Since $v$ is not a sink in $\vec{G}$ due to $\overrightarrow{vu^{+}}$, $t$ is the unique sink in $\vec{G}$. This ends the proof of ˜2.2.1.
@@
We claim that $u^{+}=t$. Suppose for contradiction that $u^{+}\neq t$. Then, since D2 is satisfied in $\vec{G}$, there exists $u^{++}\in N^{+}(u^{+})$ such that $N^{+}(u^{+})\subseteq N[u^{++}]$. Since we have assumed that $v\in N^{+}(u^{+})$, it must be that $v\in N[u^{++}]$. Suppose $v\neq u^{++}$, that is, $v\in N(u^{++})$. Then since $v$ is dominated by $u$, the node $u^{++}\in N(u)$. However, this means that the vertices $v,u,u^{+}$ and $u^{++}$ induce a $K_{4}$, hence this case is not possible. So suppose $v=u^{++}$. Suppose that $u^{+}$ has some outneighbor $w$ distinct from $v$. Then, since $v=u^{++}$, $w$ must be in $N(v)$ and $u$ is connected to $w$ as $u$ dominates $v$. In this case, the vertices $v,u,u^{+}$ and $w$ induce a $K_{4}$ so this setting is not possible and it must be that $N^{+}(u^{+})=\{u^{++}\}=\{v\}$. Therefore, it must be that $u^{+}=t$ in $\vec{G}^{\prime}$, contradicting the initial assumption.
Consider now $u^{+}=t$ in $\vec{G}^{\prime}$, implying that $v$ is the only outneighbor of $u^{+}$ in $\vec{G}$. Let $z\in N(v)$. Then, $z$ must also be neighboring to $u$ as $u$ dominates $v$. We claim that $v$ is the unique sink in $\vec{G}$. The orientation $\overrightarrow{uz}$ means that the edge $u^{+}z$ must also exist implying that $u,v,u^{+}$ and $z$ induce a $K_{4}$ which is not possible. So $uz$ must be oriented as $\overrightarrow{zu}$. From D3, $zv$ should be oriented as $\overrightarrow{zv}$. Hence every neighbor of $v$ in $\vec{G}$ must be incoming to $v$. Therefore, $v$ is a sink in $\vec{G}$ and since the unique sink $u^{+}$ of $\vec{G}^{\prime}$ is not a sink in $\vec{G}$ due to $\overrightarrow{u^{+}v}$, $v$ is the unique sink in $\vec{G}$. This ends the proof of ˜2.2.2. This ends the proof of ˜2.2. This ends the proof of ˜2. Theorem 4.1 follows from the analysis in Cases 1 and 2.

### v2

@@ 2.1 Model
3 Warm Up: Labelings for Trees and Chordal Graphs
3.1 Unique Sink in Trees
We give a simple labeling that guarantees a unique sink in the class of trees. While not new (e.g., one could easily construct a certificate following the self–stabilizing leader election algorithm of [antonoiu1996trees]), we believe it gives a good intuition behind the labelings and proofs in the following sections. The label of each node consists of the orientation of edges incident to the node. A correct configuration is defined by two conditions, T1 and T2. If both T1,T2 hold at every node of a tree $T$, then $T$ has a unique sink which we define to be the leader.
@@ 3.1 Unique Sink in Trees
In an anonymous tree $T$ where the nodes of $T$ do not have any input, any configuration $\vec{T}$ satisfying T1 and T2 contains a unique sink.
Proof 3.2.
@@
3.2 Unique Sink and Acyclic Orientation in Chordal Graphs
We generalize the labeling scheme from Section 3.1 to a set of conditions guaranteeing acyclicity and a unique sink in chordal graphs. Each node’s label is the orientation of its incident edges. We refer to the nodes of $G$ and the labeling as a configuration $\vec{G}$. A graph is chordal if every induced cycle is of length three. Every chordal graph has a simplicial vertex $v$, that is, a vertex $v$ such that $N(v)$ is a clique. We define a configuration to be correct for leader election, if for every vertex $v$ of a chordal graph, the following local conditions hold:
@@ 3.2 Unique Sink and Acyclic Orientation in Chordal Graphs
Observe that checking C3 is possible, since the view by definition contains the orientation of edges between neighboring nodes.
@@
[Existence] A correct configuration with respect to C1–C3 exists and can be constructed as follows: pick a simplicial vertex $u$, direct all edges incident to $u$ toward $N(u)$, then iteratively restrict the graph to the subgraph of remaining unoriented edges and orient edges away from simplicial vertices.
@@
Suppose now that $u$ has no outneighbor, (i.e., $u$ is a sink of $\vec{G}$). For any neighbor $v$ of $u$, any outneighbor $w$ of $v$ in $\vec{G}\setminus\{u\}$ must be a neighbor of $u$ in $\vec{G}$ from C2. Consequently, $N^{+}(v)\subseteq N[u]$. We claim that there exists a vertex $t^{\prime}\in N(u)$ that has no outneighbor in $N(u)$. Suppose otherwise. Then, every $v\in N(u)$ has an outneighbor in $N(u)$ and $\vec{G}\setminus\{u\}$ contains a cycle, which is a contradiction by Lemma 3.3. Since $N(u)$ is a complete graph, this $t^{\prime}$ is unique. Furthermore, as we have already shown that $t^{\prime}$ has no outneighbor outside of $N[u]$, $t^{\prime}$ is the unique sink $t$ of $\vec{G}\setminus\{u\}$. Since $^{\prime}=t\in N^{-}(u), tt$ is not a sink in $\vec{G}$. Therefore, $u$ is the unique sink of $\vec{G}$.
Theorem 3.7 directly follows from Section˜3.2 and Lemmas˜3.3 and 3.5.
@@
Observe that to construct a spanning tree in a chordal graph satisfying C1–C3, it is enough that each non-root node picks an outgoing edge.
@@ 4 Dismantlable graphs
[…]
@@ 4.1 Leader election in $K_{4}$-free dismantlable graphs
Here we only consider $K_{4}$-free dismantlable graphs. As in Section 3.2, the label at each node consists of an orientation on the edges incident to the node and we write $\vec{G}$ to denote a configuration. We define a locally correct configuration to be one that satisfies D1–D3 for every node $v$ of $G$. We show that if all rules are satisfied, $G$ contains a unique sink.
@@
[Existence] Any dismantlable graph admits an acyclic orientation satisfying D1–D3. {claimproof} Let $v_{1},v_{2}\ldots,v_{n}$ be the dismantling order of the nodes of some $K_{4}$-free dismantlable graph $G$. Orient each edge between two nodes $v_{i}$ and $v_{j}$ as $\overrightarrow{v_{i}v_{j}}$ if and only if $i>j$ in the dismantling order. An example of this orientation for the dismantiling order $(i,h,g,f,e,d,c,b,a)$ of the graph in Figure 3 is given in Figure 3. This orientation satisfies D1 and D3. For D2, consider any node $v_{i}$. If $i=1$, then $N^{+}(v)=\emptyset$, and D2 is immediately satisfied for $v$. Suppose now that $i>1$. Since $v_{1},v_{2},\ldots,v_{n}$ is a dismantling order, there exists $j<i$ such that $N[v_{i}]\cap\{v_{1},\ldots v_{i}\}\subseteq N[v_{j}]$. By the definition of the orientation of $G$, $N^{+}(v_{i})=N(v_{i})\cap\{v_{1},\ldots v_{i}\}$ and thus $N^{+}(v_{i})\subseteq N[v_{j}]$, i.e., D2 holds for $v_{i}$ with $v_{i}^{+}=v_{j}$.
Notice that in the case of $K_{4}$-free dismantlable graphs, one cannot hope to have a property similar to Lemma˜3.3. As in [chalopin2024stabilising], one can construct configurations satisfying D1-D3 that contain directed cycles (e.g., Figure˜3) and are thus not obtained from a dismantling order.
Figure 3: (a) An example of a $K_{4}$-free dismantlable graph. (b) A $K_{4}$-free dismantlable graph and an acyclic labeling that satisfies D1–D3. (c) A $K_{4}$-free dismantlable graph and a labeling that satisfies D1–D3 but is not acyclic. The edges forming the cycle are marked in orange.
@@
In the remainder of the proof, we distinguish different cases. In each case, we first show that D2 is satisfied in $\vec{G}^{\prime}$, then we apply the induction hypothesis on $\vec{G}^{\prime}$ to establish that $\vec{G}^{\prime}$ contains a unique sink, and finally, we deduce that $\vec{G}$ also contains a unique sink. We distinguish two cases, depending on whether $uv$ is directed as $\overrightarrow{uv}$ or $\overrightarrow{vu}$.
@@
We first show that D2 is satisfied in $\vec{G}^{\prime}$. Consider a vertex $w\neq v$ such that $w\in N^{-}(v)$. If $w\neq u$, since $u$ dominates $v$, $u$ must be a neighbor of $w$. Given that $N^{+}(u)\cap N^{-}(v)=\emptyset$, it follows that $w\in N^{-}(u)$. By Lemma˜4.3, we can assume $w^{+}\neq v$, thus D2 is satisfied for $w$ in $\vec{G}^{\prime}$. We now show that D2 holds for $u$. Since $N^{+}(u)\subseteq N[v]$ and $N^{+}(u)\cap N^{-}(v)=\emptyset$, we have $N^{+}(u)\subseteq N^{+}(v)\cup\{v\}$. Given that $N^{+}(v)\subseteq N(v)\subseteq N[u]$ and $v\in N^{+}(u)$, by D3, we obtain $N^{+}(v)\subseteq N^{+}(u)$. Therefore, $N^{+}(u)=N^{+}(v)\cup\{v\}$. Hence, either $N^{+}(v)=\emptyset$ and $N^{+}(u)=\{v\}$, or there exists $v^{+}\in N^{+}(v)$ such that $N^{+}(v)\subseteq N[v^{+}]$, implying $N^{+}(u)=N^{+}(v)\cup\{v\}\subseteq N[v^{+}]$. In either case, D2 is satisfied for $u$ in $\vec{G}^{\prime}$. Observe also that as $N[v]\subseteq N[u]$ and as $N^{+}(u)=N^{+}[v]$, necessarily we have $N^{-}(v)\subseteq N^{-}[u]$.
By induction hypothesis $\vec{G}^{\prime}$ contains a unique sink $t$. Notice that as $N^{+}(u)=N^{+}[v]$, $v$ is a sink of $\vec{G}$ if and only if $u$ is a sink of $\vec{G}^{\prime}$. Therefore, if $t=u$, then $t$ is not a sink in $\vec{G}$ but $v$ is, and consequently $v$ is the unique sink of $\vec{G}$. Suppose now that $t\neq u$. Then $v$ is not a sink in $\vec{G}$ as $N^{+}[v]=N^{+}(u)$ is non-empty. Moreover, since $N^{-}(v)\subseteq N^{-}[u]$ and since $t\notin N^{-}(u)$, $v\notin N^{+}(t)$ and $t$ is also a sink in $\vec{G}$. Thus, $t$ is the unique sink in $\vec{G}$. In either case, $\vec{G}$ contains a unique sink concluding the proof of ˜2.1.1.
@@
We now show that D2 holds in $\vec{G}^{*}$. For any $\notin(N^{+}(u)\cap N^{-}(v))\cup\{u\}, xx$ has the same set of outneighbors in $\vec{G}$ and in $\vec{G}^{*}$, and thus D2 holds for $x$ in $\vec{G}^{*}$ as it holds in $\vec{G}$. Since $N^{+}(u)\subseteq N(v)$ in $\vec{G}$, since $N^{+}(u)$ in $\vec{G}^{*}$ is a subset of $N^{+}(u)$ in $\vec{G}$, and since $v\in N^{+}(u)$ in $\vec{G}^{*}$, D2 also holds for $u$ in $\vec{G}^{*}$ (with $u^{+}=v$). Consider now $w\in(N^{+}(u)\cap N^{-}(v))$ and let $w^{+}\in N^{+}(w)$ such that $N^{+}(w)\subseteq N[w^{+}]$ in $\vec{G}^{\prime}$. As $u\notin N^{+}(w)$ in $\vec{G}$, $w^{+}\neq u$. Since $v\in N^{+}(w)$, $w^{+}\in N[v]\setminus\{u\}\subseteq N(u)$. Hence, $w^{+}\in N(u)$, implying $u\in N[w^{+}]$ and thus in $\vec{G}^{*}$ we also have $N^{+}(w)\subseteq N[w^{+}]$. Therefore, D2 is satisfied at $w$ in $\vec{G}^{*}$. So D2 holds in $\vec{G}^{*}$.
@@
We first show that D2 is satisfied in $\vec{G}^{\prime}$. More precisely, we show that for each $w\neq v$, either $N^{+}(w)\setminus\{v\}=\emptyset$, or there exists $w^{+}\neq v$ such that $N^{+}(w)\subseteq N[w^{+}]$ in $\vec{G}$ (and thus in $\vec{G^{\prime}}$). Note that if $N^{+}(w)=\emptyset$, the property holds. Otherwise, there exists $w^{+}$ such that $N^{+}(w)\subseteq N[w^{+}]$ in $G$. If $w^{+}\neq v$ or if $N^{+}(w)=\{v\}$, we are done. Suppose now that $w^{+}=v$. Since $w\in N(v)$, we have $w\in N(u)$, and if $u\in N^{+}(w)$, by Lemma˜4.3, we can assume that $w^{+}=u$ and we are also done. Suppose now that $w^{+}=v$, that $u\in N^{-}(w)$, and that there exists $z\in N^{+}(w)\setminus\{v\}$. Then $z\in N(v)\subseteq N[u]$ and $z\neq u$ as $w\in N^{-}(z)\cap N^{+}(u)$. But then $v,u,w,z$ induce a $K_{4}$, which is a contradiction. This shows that D2 holds for all nodes in $\vec{G}^{\prime}$.
By induction hypothesis, $\vec{G}^{\prime}$ contains a unique sink $t$. For any vertex $w$ of $\vec{G^{\prime}}$ distinct from $t$, we have just shown that we can assume that $w^{+}\neq v$. We consider two cases, based on whether $t=u^{+}$ or not. Note that in both cases, $v\in N^{+}(u)\subseteq N[u^{+}]$ and thus $v$ is a neighbor of $u^{+}$.
@@
Since $u^{+}$ is the unique sink of $\vec{G}^{\prime}$, if $v\in N^{-}(u^{+})$, then $u^{+}$ is the unique sink of $\vec{G}$. Suppose now that $v\in N^{+}(u^{+})$ and observe that $v$ is the only potential sink in $\vec{G}$. We claim that $v$ is indeed the unique sink in $\vec{G}$. Suppose this is not true and that there exists $z\in N^{+}(v)$. Note that $z\notin\{u,u^{+}\}$ as $v\in N^{-}(z)\cap N^{+}(u)\cap N^{+}(u^{+})$. Since $N(v)\subseteq N[u]$, $z$ is also a neighbor of $u$. From D3 applied to the triangle uvzuvz, $z\in N^{+}(u)$ and thus $z\in N[u^{+}]$. Consequently, $v,u,u^{+}$ and $z$ induce a $K_{4}$ which is impossible. This ends the proof of ˜2.2.1.
@@
If $t$ is not a sink in $\vec{G}$, $t\in N^{-}(v)$ and thus, $t\in N[u]$. Since $t$ is a sink in $\vec{G}^{\prime}$, $t\neq u$ as $u$ has an outneighbor $u^{+}$ in $\vec{G}^{\prime}$ and $t\in N^{+}(u)$. Since $v,t\in N^{+}(u)\subseteq N[u^{+}]$ and since $u,v,u^{+},t$ are four different vertices, they induce a $K_{4}$ in $G$, which is a contradiction. Consequently, $t$ is a sink in $\vec{G}$.
Suppose now that $t$ is not the unique sink of $\vec{G}$, i.e., that $v$ is also a sink of $\vec{G}$. This implies that $u,u^{+}\in N^{-}(v)$. Since $u^{+}$ is not a sink in $\vec{G^{\prime}}$, it is not a sink in $\vec{G}$ either and thus there exists $u^{++}\in V(G)$ such that $N^{+}(u^{+})\subseteq N[u^{++}]$. Since we can assume that $u^{++}\neq v$ and since $v\in N^{+}(u^{+})\subseteq N[u^{++}]$, we have $u^{++}\in N(v)\subseteq N[u]$. Note that $u^{++}\neq u$ since $u^{+}\in N^{+}(u)\cap N^{-}(u^{++})$ Therefore $v,u,u^{+},u^{++}$ induce a $K_{4}$, which is a contradiction. This establishes that $t$ is the unique sink of $\vec{G}$. This ends the proof of ˜2.2.2. This ends the proof of ˜2.2. This ends the proof of ˜2. Theorem 4.1 follows from the analysis in Cases 1 and 2.